how many grams of My need to react to produce 210. mL of H2 collected over water at 25.0°C and a barometric pressure of 743 Torr. The vapor pressure of water at 25.0°C is 24 torr.

Mg (s) + 2HCl => MgCl2 (aq) + H2 (g)
Answer #1

Mg (s) + 2Hd lol Myd, (aq) + H₂G) pressure of H₂= 743-24 torr - 719 torr 719 atm 760 = 0.946 atm Temperature: 25°C= 298k mole